Write the slope intercept equation for the line that passes through (-2,3) and is perpendicular to 2x-9y=-12
Answer by mananth(16946)

2 x -9 y = -12
Find the slope of this line

-9 y = -2 x -12
Divide by -9
y = 2/ 9 x + 1 1/3
Compare this equation with y=mx+b
slope, m=2/9

The slope of a line perpendicular to the above line will be the negative reciprocal -4.5
m1*m2=-1
m= -9/2 ,point (-2,3)
Find b by plugging the values of m & the point in
y=mx+b
3 = 9.00 +b
b=-6
m=-9/2
The required equation is y=-(9/ 2)x-6
m.ananth@hotmail.ca